# Galley: Modern Query Optimization for Sparse Tensor Programs

### v3

@@ 10. Conclusion & Limitations
In this paper, we presented Galley, our system for declarative sparse tensor programming. We described and then demonstrated how it optimizes high-level program structure with its logical optimizer and how it lowers that program to an efficient implementation with its physical optimizer. These decisions are guided by sparsity estimates of intermediate expressions, and we show that these estimates can be computed for arbitrary tensor algebra programs by implementing a minimal 5-function interface. Lastly, we present an experimental evaluation of this work on ML algorithms over structured feature data, sub-graph counting, and breadth-first search.
@@
There are a few optimizations which we are excited to bring to Galley in the future. Currently, it does not support complex loop structures (e.g. a single outer FOR loop which wraps multiple inner FOR loops) or parallelism. Both of these areas can benefit from cost-based optimization, and we’re excited to explore them in future work. Similarly, we do not consider memory constraints during the optimization process, but our focus on cardinality bounds provides an exciting avenue for approaching this in the future.

### v4

@@ 9. Conclusion and Limitations
This paper introduced Galley, a system for efficient declarative sparse tensor programming. We describe and then demonstrate how it optimizes high-level program structure with its logical optimizer and how it lowers that program to an efficient implementation with its physical optimizer. Guiding these decisions are sparsity estimates of intermediate expressions, and we show that these estimates can be computed for arbitrary tensor algebra programs by implementing a minimal 5-function interface. Finally, we present results from our evaluation of experiments with Galley on ML algorithms over structured feature data, sub-graph counting, and breadth-first search.
@@
We are excited to enrich Galley further with new optimizations in the future. Currently, Galley lacks support for complex loop structures (e.g., a single outer FOR loop that wraps multiple inner FOR loops) and parallelism. Both areas could benefit from cost-based optimization. Similarly, Galley does not consider hard memory constraints during optimization, but our use of cardinality bound methods provides an avenue for addressing this in future work.
Acknowledgements.
This work was supported by NSF SHF 2312195, NSF IIS 2314527, and DARPA SBIR: HR001123C0139.